# Sequential Diversification with Provable Guarantees

### v2

@@ C.2. The case of an arbitrary ranking
$\mathcal{S}_{+}(O)$ $=\sum_{i=1}^{n-1}p_{O_{i+1}}d(\pi(i+1),O_{i})$
$\geq\sum_{i=1}^{n-1}\left(c-\Theta\left(\frac{1}{n}\right)\right)d(\pi(i+1),O_{i})$
$=\left(c-\Theta\left(\frac{1}{n}\right)\right)\sum_{i=1}^{n-1}d(\pi(i+1),O_{i})$
@@ Appendix D Omitted proofs from Section 7
$\mathcal{H}(O)=\sum_{i=1}^{n-1}W_{O_{i}}d(\pi(i),\pi(i+1)).$
Let $\pi^{o}$ be the optimal ordering for the MaxOHP. Let $O_{\pi^{o}}=(\pi^{o}(i))_{i=1}^{n}$ and write $O^{o}=O_{\pi^{o}}$. For the optimal ordering $O^{o}$, let $\mathcal{H}(O^{o})=\sum_{i=1}^{n-1}W_{O_{i}^{o}}d(\pi^{o}(i),\pi^{o}(i+1))$.
@@
$\mathcal{H}(O)$ $=\sum_{i=1}^{n-1}W_{O_{i}}d(\pi(i),\pi(i+1))$
$\geq\sum_{i=1}^{\tau-1}W_{O_{i}}d(\pi(i),\pi(i+1))$
$\geq\sum_{i=1}^{\tau-1}\sum_{j=i+1}^{\tau}p_{O_{j}}d(\pi(i),\pi(i+1)).$
@@
$\mathcal{H}$ $(O^{o})=\sum_{i=1}^{n-1}W_{O_{i}^{o}}d(\pi^{o}(i),\pi^{o}(i+1))$
$=\sum_{i=1}^{n-1}\sum_{j=i+1}^{n}p_{O^{o}_{j}}d(\pi^{o}(i),\pi^{o}(i+1))$
$=\sum_{t=1}^{T}\sum_{i=(t-1)(\tau-1)+1}^{t(\tau-1)}\sum_{j=i+1}^{n}p_{O^{o}_{j}}d(\pi^{o}(i),\pi^{o}(i+1))$
$=\sum_{t=1}^{T}\sum_{i=(t-1)(\tau-1)+1}^{t(\tau-1)}\left(\sum_{j=i+1}^{t\cdot\tau}p_{O^{o}_{j}}+\sum_{j=t\cdot\tau+1}^{n}p_{O^{o}_{j}}\right)d(\pi^{o}(i),\pi^{o}(i+1))$
@@
(D.1) $\sum_{i=(t-1)(\tau-1)+1}^{t(\tau-1)}\sum_{j=i+1}^{t\cdot\tau}p_{O^{o}_{j}}d(\pi^{o}(i),\pi^{o}(i+1))$
(D.2) $\sum_{i=(t-1)(\tau-1)+1}^{t(\tau-1)}\sum_{j=t\cdot\tau+1}^{n}p_{O^{o}_{j}}d(\pi^{o}(i),\pi^{o}(i+1))$
@@
(D.4) $\sum_{i=(t-1)(\tau-1)+1}^{t(\tau-1)}$ $\sum_{j=t\cdot\tau+1}^{n}p_{O^{o}_{j}}d(\pi^{o}(i),\pi^{o}(i+1))$
$\stackrel{{\scriptstyle(b)}}{{\leq}}\sum_{i=(t-1)(\tau-1)+1}^{t(\tau-1)}\sum_{j=t\cdot\tau+1}^{n}b^{j}d(\pi^{o}(i),\pi^{o}(i+1))$
$\leq\sum_{i=(t-1)(\tau-1)+1}^{t(\tau-1)}\frac{b^{t\cdot\tau+1}}{1-b}d(\pi^{o}(i),\pi^{o}(i+1))$
@@
where $d_{\max}\coloneqq\max_{u,v\in U}d(u,v)$. Inequality $(a)$ holds by the definition of $\ell_{\tau}$ and our algorithm, and inequality $(b)$ holds because $p_{O^{o}_{j}}\leq b^{j}$.
@@
(D.6) $\ell_{\tau}$ $\stackrel{{\scriptstyle(a)}}{{\geq}}\ell_{2}\stackrel{{\scriptstyle(b)}}{{\geq}}p_{\hat{O_{2}}}d(\hat{\pi}(1),\hat{\pi}(2))\geq a^{2}d_{\max}.$
@@
[…]
Random EXPLORE DUM MSD MMR DPP B2I B3I B4I
Coat 0.039 $\pm$ 0.020 0.105 $\pm$ 0.041 0.102 $\pm$ 0.038 0.095 $\pm$ 0.039 0.105 $\pm$ 0.041 0.105 $\pm$ 0.041 0.109 $\pm$ 0.042 0.106 $\pm$ 0.038 0.104* $\pm$ 0.037*
KuaiRec 0.022 $\pm$ 0.006 0.050 $\pm$ 0.000 0.033 $\pm$ 0.008 0.042 $\pm$ 0.006 0.051 $\pm$ 0.010 0.044 $\pm$ 0.006 0.058 $\pm$ 0.010 0.058 $\pm$ 0.009* 0.058 $\pm$ 0.009*
Netflix 0.063 $\pm$ 0.026 0.152 $\pm$ 0.026 0.145 $\pm$ 0.026 0.161 $\pm$ 0.024 0.162 $\pm$ 0.024 0.162 $\pm$ 0.024 0.163 $\pm$ 0.024 0.160 $\pm$ 0.025 0.159* $\pm$ 0.025*
Movielens 0.059 $\pm$ 0.028 0.173 $\pm$ 0.013 0.169 $\pm$ 0.018 0.183 $\pm$ 0.003 0.183 $\pm$ 0.003 0.183 $\pm$ 0.003 0.183 $\pm$ 0.003 0.183 $\pm$ 0.003* 0.183 $\pm$ 0.003*
Yahoo 0.016 $\pm$ 0.027 0.133 $\pm$ 0.044 0.148 $\pm$ 0.044 0.148 $\pm$ 0.044 0.148 $\pm$ 0.044 0.148 $\pm$ 0.044 0.148 $\pm$ 0.044 0.140 $\pm$ 0.045* 0.139 $\pm$ 0.045*
LETOR 0.008 ± 0.008 0.026 ± 0.021 0.043 ± 0.028 0.039 ± 0.026 0.040 ± 0.026 0.040 ± 0.026 0.043 ± 0.028 0.042 ± 0.027 0.042 ± 0.027
LTRC 0.019 ± 0.011 0.041 ± 0.018 0.054 ± 0.020 0.051 ± 0.020 0.051 ± 0.019 0.051 ± 0.019 0.055 ± 0.020 0.054 ± 0.020 0.054 ± 0.020
[…]
Random EXPLORE DUM MSD MMR DPP B2I B3I B4I
Coat 9.703 $\pm$ 3.704 23.044 $\pm$ 12.390 22.251 $\pm$ 11.328 18.586 $\pm$ 8.481 24.151 $\pm$ 13.421 24.213 $\pm$ 13.748 23.030 $\pm$ 12.385 22.839 $\pm$ 12.428* 22.826 $\pm$ 12.415*
KuaiRec 5.737 $\pm$ 1.094 10.668 $\pm$ 1.874 11.599 $\pm$ 3.289 9.119 $\pm$ 1.090 14.654 $\pm$ 3.874 12.565 $\pm$ 4.258 10.920 $\pm$ 1.728 10.906 $\pm$ 1.714* 10.968 $\pm$ 1.726*
Netflix 16.901 $\pm$ 5.588 47.316 $\pm$ 14.231 39.696 $\pm$ 9.856 35.729 $\pm$ 10.010 48.595 $\pm$ 14.424 50.530 $\pm$ 15.407 47.224 $\pm$ 14.648 47.053 $\pm$ 14.294* 47.172 $\pm$ 14.266*
Movielens 14.677 $\pm$ 4.644 65.410 $\pm$ 6.715 48.412 $\pm$ 4.246 63.242 $\pm$ 8.458 69.341 $\pm$ 7.604 71.362 $\pm$ 7.006 71.303 $\pm$ 7.748 71.305 $\pm$ 7.738* 71.317 $\pm$ 7.703*
Yahoo 4.014 $\pm$ 3.259 40.505 $\pm$ 19.773 38.251 $\pm$ 18.866 38.520 $\pm$ 19.221 39.134 $\pm$ 19.191 38.377 $\pm$ 19.021 41.677 $\pm$ 21.090 41.924 $\pm$ 20.996* 41.969 $\pm$ 21.010*
LETOR 2.784 ± 1.638 5.825 ± 3.450 5.317 ± 2.755 6.140 ± 2.825 6.523 ± 3.803 6.025 ± 3.637 6.034 ± 3.006 6.086 ± 3.033 6.165 ± 3.048
LTRC 4.757 ± 1.339 7.026 ± 2.161 6.332 ± 1.691 7.088 ± 1.958 7.385 ± 2.352 7.204 ± 2.164 6.972 ± 1.971 7.005 ± 1.990 7.058 ± 2.012
[…]
Random EXPLORE DUM MSD MMR DPP B2I B3I B4I
Coat 0.555±0.722 7.665±10.628 8.409±10.708 9.660±13.374 9.661±13.608 9.559±13.534 9.075±12.593 9.119±12.67*7 9.100±12.686*
KuaiRec 0.255±0.191 1.140±0.554 4.808±3.282 6.388±4.981 6.768±5.221 6.739±5.140 3.726±2.575 3.745±2.570* 3.834±2.617*
Netflix 1.704±1.822 27.252±19.761 22.197±12.660 32.270±22.224 32.703±22.895 33.157±23.552 30.790±21.842 30.848±21.822* 30.915±21.813*
Movielens 1.157±1.283 61.736±9.103 37.570±5.342 59.667±13.080 68.657±9.789 66.995±13.290 69.476±11.786 69.478±11.780* 69.484±11.759*
Yahoo 0.394±0.927 26.104±24.181 21.018±20.291 22.055±20.980 26.303±23.451 21.039±20.324 27.567±26.163 27.709±26.100* 27.746±26.105*
LETOR 0.127 ± 0.137 0.436 ± 0.419 0.583 ± 0.478 0.595 ± 0.491 0.612 ± 0.521 0.592 ± 0.510 0.610 ± 0.515 0.608 ± 0.516 0.615 ± 0.520
LTRC 0.189 ± 0.216 0.702 ± 0.640 1.050 ± 0.865 1.096 ± 0.914 1.136 ± 0.996 1.136 ± 1.000 1.091 ± 0.899 1.099 ± 0.911 1.112 ± 0.928
Appendix E Experiments and results
In this section, we present the omitted dataset information, baseline description, the experimental setting, and more experimental results.
E.1. Datasets
Coat: [^2] https://www.cs.cornell.edu/~schnabts/mnar/ Coat ratings in the range $[1,5]$, accompanied by categorical features, such as gender, jacket type, and color.
Netflix: [^3] https://www.kaggle.com/datasets/rishitjavia/netflix-movie-rating-dataset, [^4] https://github.com/tommasocarraro/netflix-prize-with-genres Movie ratings in range $[1,5]$ with genre information. We sample $5$ K movies and exclude users with less than $20$ interactions.
Movielens: [^5] https://grouplens.org/datasets/movielens/1m/ Movie ratings in range $[1,5]$ with movie genres.
Yahoo: [^6] https://webscope.sandbox.yahoo.com/catalog.php?datatype=i&did=67 Song ratings in range $[1,5]$ with song genres. We sample $3$ K songs and exclude users with less than $20$ interactions.
LETOR: [^7] https://www.microsoft.com/en-us/research/project/letor-learning-rank-information-retrieval/letor-4-0/ Web-search dataset with relevance scores (in $\{0,1,2\}$) on queries-document pairs and document feature vectors.
LTRC: [^8] https://webscope.sandbox.yahoo.com/catalog.php?datatype=c Web-search dataset with relevance scores (in $\{0,1,2,3,4\}$) on queries-document pairs and document feature vectors.
E.2. Omitted Baseline
@@
$[p_{i}^{-\alpha}+d(i,R)^{-\alpha}-1]^{-1/\alpha}.$
@@
E.3. More experimental results
Experimental environment All experiments are conducted on a dual-socket server, each socket housing an Intel Xeon Gold 6326 CPU at 2.90GHz, with 128 cores and 256 threads.
[…]
The EXPLORE algorithm, despite designed to maximize the sequential diversity, generally under-performs compared to $\textsc{B}\tau\textsc{I}$ in terms of sequential diversity value. This may be due to the different settings, as EXPLORE is designed for scenarios where users interact with multiple lists, accepting only one item from each, whereas $\textsc{B}\tau\textsc{I}$ allows users to sequentially accept items from a single list.
[…]
Metrics Datasets Random EXPLORE DUM MSD MMR DPP B2I B3I B4I
[…]
KuaiRec 0.040 ± 0.012 0.080±0.017 0.101 ± 0.013 0.085 ± 0.010 0.091 ± 0.014 0.085 ± 0.010 0.070 ± 0.023 0.090 ± 0.011* 0.091 ± 0.011*
Netflix 0.054 ± 0.020 0.095±0.017 0.105 ± 0.012 0.105 ± 0.012 0.105 ± 0.012 0.105 ± 0.012 0.103 ± 0.012 0.104 ± 0.012* 0.104 ± 0.012*
Movielens 0.054 ± 0.025 0.107±0.010 0.112 ± 0.001 0.112 ± 0.001 0.112 ± 0.001 0.112 ± 0.001 0.112 ± 0.001 0.112 ± 0.001* 0.112 ± 0.001*
Yahoo 0.044 ± 0.026 0.090±0.027 0.099 ± 0.022 0.099 ± 0.022 0.099 ± 0.022 0.099 ± 0.022 0.095 ± 0.025 0.099 ± 0.022* 0.099 ± 0.022*
LETOR 0.022 ± 0.022 0.030±0.028 0.050 ± 0.038 0.065 ± 0.036 0.065 ± 0.036 0.066 ± 0.036 0.050 ± 0.038 0.063 ± 0.035 0.063 ± 0.035
LTRC 0.033 ± 0.019 0.042±0.024 0.056 ± 0.024 0.070 ± 0.026 0.070 ± 0.026 0.069 ± 0.025 0.056 ± 0.024 0.064 ± 0.025 0.064 ± 0.025
[…]
Netflix 0.005 ± 0.015 0.014±0.027 0.019 ± 0.033 0.019 ± 0.032 0.019 ± 0.032 0.019 ± 0.032 0.018 ± 0.031 0.017 ± 0.030* 0.017 ± 0.030*
Movielens 0.006 ± 0.018 0.018±0.029 0.019 ± 0.032 0.013 ± 0.026 0.013 ± 0.026 0.013 ± 0.026 0.013 ± 0.026 0.013 ± 0.026* 0.013 ± 0.026*
Yahoo 0.003 ± 0.012 0.020±0.021 0.021 ± 0.017 0.019 ± 0.015 0.019 ± 0.015 0.019 ± 0.015 0.027 ± 0.023 0.017 ± 0.014* 0.017 ± 0.014*
[…]
Metrics Datasets Random EXPLORE DUM MSD MMR DPP B2I B3I B4I
[…]
KuaiRec 1.565 ± 0.162 2.034±0.174 2.789 ± 0.308 1.772 ± 0.104 2.416 ± 0.443 2.905 ± 0.350 1.772 ± 0.145 1.887 ± 0.130* 1.891 ± 0.131*
Netflix 1.889 ± 0.280 2.881±0.365 2.880 ± 0.308 2.681 ± 0.290 3.003 ± 0.350 2.993 ± 0.349 2.841 ± 0.338 2.855 ± 0.332* 2.865 ± 0.330*
Movielens 1.782 ± 0.287 3.280±0.101 3.141 ± 0.069 3.213 ± 0.109 3.282 ± 0.097 3.327 ± 0.075 3.286 ± 0.097 3.287 ± 0.096* 3.288 ± 0.093*
Yahoo 1.633 ± 0.405 2.704±0.605 2.636 ± 0.556 2.637 ± 0.556 2.644 ± 0.557 2.637 ± 0.556 2.653 ± 0.578 2.698 ± 0.553* 2.703 ± 0.553*
LETOR 1.228 ± 0.281 1.578±0.443 1.402 ± 0.355 1.617 ± 0.411 1.748 ± 0.564 1.834 ± 0.575 1.400 ± 0.341 1.523 ± 0.365 1.545 ± 0.373
LTRC 1.438 ± 0.235 1.659±0.278 1.562 ± 0.253 1.744 ± 0.281 1.833 ± 0.341 1.814 ± 0.310 1.567 ± 0.253 1.615 ± 0.259 1.637 ± 0.263
[…]
Netflix 0.280 ± 0.386 0.965±0.892 1.187 ± 0.900 1.336 ± 0.852 0.830 ± 0.848 1.042 ± 0.952 1.290 ± 0.936 1.287 ± 0.942* 1.285 ± 0.949*
Movielens 0.289 ± 0.422 1.253±1.078 1.073 ± 0.984 1.152 ± 1.053 1.201 ± 1.151 1.183 ± 1.146 1.209 ± 1.097 1.209 ± 1.097* 1.210 ± 1.097*
Yahoo 0.229 ± 0.300 2.754±1.380 3.858 ± 1.547 3.857 ± 1.546 3.833 ± 1.537 3.852 ± 1.540 3.859 ± 1.563 3.675 ± 1.659* 3.643 ± 1.668*
[…]
Metrics Datasets Random EXPLORE DUM MSD MMR DPP B2I B3I B4I
[…]
KuaiRec 0.299 ± 0.181 0.925±0.229 2.100 ± 0.658 2.259 ± 0.760 2.201 ± 0.743 2.161 ± 0.717 1.271 ± 0.436 1.399 ± 0.415* 1.453 ± 0.430*
Netflix 0.581 ± 0.341 2.221±0.772 2.335 ± 0.674 2.492 ± 0.777 2.447 ± 0.765 2.481 ± 0.773 2.347 ± 0.749 2.368 ± 0.739* 2.378 ± 0.738*
Movielens 0.525 ± 0.369 3.251±0.191 3.005 ± 0.129 3.347 ± 0.083 3.294 ± 0.119 3.281 ± 0.163 3.274 ± 0.157 3.274 ± 0.156* 3.275 ± 0.152*
Yahoo 0.407 ± 0.423 2.074±1.064 1.991 ± 0.947 2.020 ± 0.944 2.278 ± 0.995 1.991 ± 0.947 2.051 ± 1.031 2.131 ± 0.984* 2.137 ± 0.986*
LETOR 0.199 ± 0.150 0.304±0.219 0.377 ± 0.258 0.514 ± 0.282 0.513 ± 0.300 0.523 ± 0.291 0.376 ± 0.260 0.465 ± 0.253 0.472 ± 0.260
LTRC 0.233 ± 0.220 0.433±0.359 0.667 ± 0.407 0.835 ± 0.502 0.841 ± 0.518 0.855 ± 0.525 0.666 ± 0.408 0.737 ± 0.433 0.751 ± 0.443
[…]
Netflix 0.068 ± 0.164 0.715±0.790 0.836 ± 0.801 0.680 ± 0.774 0.869 ± 0.893 0.816 ± 0.889 0.871 ± 0.876 0.873 ± 0.881* 0.876 ± 0.886*
Movielens 0.055 ± 0.156 1.265±1.125 1.030 ± 0.931 0.744 ± 0.998 1.179 ± 1.134 1.164 ± 1.139 1.223 ± 1.099 1.223 ± 1.099* 1.223 ± 1.099*
Yahoo 0.051 ± 0.143 1.842±1.588 2.461 ± 1.974 2.452 ± 1.975 2.143 ± 1.987 2.448 ± 1.954 2.505 ± 2.043 2.439 ± 2.076* 2.426 ± 2.078*
[…]
Random EXPLORE DUM MSD MMR DPP B2I B3I B4I
Coat 0.83 2.81 0.97 1.08 1.24 0.96 1.41 4.70* 53.03*
KuaiRec 21.00 134.82 25.40 43.92 58.37 106.47 198.61 55.27* 283.62*
Netflix 10.76 180.60 14.51 14.71 14.51 50.67 181.57 72.79* 866.86*
Movielens 40.12 787.52 68.91 135.89 87.50 561.19 929.23 154.21* 1121.35*
Yahoo 53.78 1922.62 132.07 252.97 125.09 1316.29 2089.38 388.58* 3807.37*
LETOR 0.17 9.90 2.71 1.94 1.74 2.12 0.93 10.05 873.70
LTRC 0.45 28.30 6.52 4.98 4.97 5.95 2.24 42.11 3584.67

### v3

@@ C.2. The Case of an Arbitrary Ranking
$\mathcal{S}_{+}(O)$ $=\sum_{i=1}^{n-1}p_{O_{i+1}}d\left(\pi(i+1),O_{i}\right)$
$\geq\sum_{i=1}^{n-1}\left(c-\Theta\left(\frac{1}{n}\right)\right)d\left(\pi(i+1),O_{i}\right)$
$=\left(c-\Theta\left(\frac{1}{n}\right)\right)\sum_{i=1}^{n-1}d\left(\pi(i+1),O_{i}\right)$
@@ Appendix D Omitted Proofs from Section 7
$\mathcal{H}(O)=\sum_{i=1}^{n-1}W_{O_{i}}d\left(\pi(i),\pi(i+1)\right).$
Let $\pi^{o}$ be the optimal ordering for the MaxOHP. Let $O_{\pi^{o}}=(\pi^{o}(i))_{i=1}^{n}$ and write $O^{o}=O_{\pi^{o}}$. For the optimal ordering $O^{o}$, let $\mathcal{H}(O^{o})=\sum_{i=1}^{n-1}W_{O_{i}^{o}}d\left(\pi^{o}(i),\pi^{o}(i+1)\right)$.
@@
$\mathcal{H}(O)$ $=\sum_{i=1}^{n-1}W_{O_{i}}d\left(\pi(i),\pi(i+1)\right)$
$\geq\sum_{i=1}^{\tau-1}W_{O_{i}}d\left(\pi(i),\pi(i+1)\right)$
$\geq\sum_{i=1}^{\tau-1}\sum_{j=i+1}^{\tau}p_{O_{j}}d\left(\pi(i),\pi(i+1)\right).$
@@
$\mathcal{H}$ $(O^{o})=\sum_{i=1}^{n-1}W_{O_{i}^{o}}d\left(\pi^{o}(i),\pi^{o}(i+1)\right)$
$=\sum_{i=1}^{n-1}\sum_{j=i+1}^{n}p_{O^{o}_{j}}d\left(\pi^{o}(i),\pi^{o}(i+1)\right)$
$=\sum_{t=1}^{T}\sum_{i=(t-1)(\tau-1)+1}^{t(\tau-1)}\sum_{j=i+1}^{n}p_{O^{o}_{j}}d\left(\pi^{o}(i),\pi^{o}(i+1)\right)$
$=\sum_{t=1}^{T}\sum_{i=(t-1)(\tau-1)+1}^{t(\tau-1)}\left(\sum_{j=i+1}^{t\cdot\tau}p_{O^{o}_{j}}+\sum_{j=t\cdot\tau+1}^{n}p_{O^{o}_{j}}\right)d\left(\pi^{o}(i),\pi^{o}(i+1)\right)$
@@
(D.1) $\sum_{i=(t-1)(\tau-1)+1}^{t(\tau-1)}\sum_{j=i+1}^{t\cdot\tau}p_{O^{o}_{j}}d\left(\pi^{o}(i),\pi^{o}(i+1)\right)$
(D.2) $\sum_{i=(t-1)(\tau-1)+1}^{t(\tau-1)}\sum_{j=t\cdot\tau+1}^{n}p_{O^{o}_{j}}d\left(\pi^{o}(i),\pi^{o}(i+1)\right)$
@@
(D.4) $\sum_{i=(t-1)(\tau-1)+1}^{t(\tau-1)}$ $\sum_{j=t\cdot\tau+1}^{n}p_{O^{o}_{j}}d\left(\pi^{o}(i),\pi^{o}(i+1)\right)$
$\stackrel{{\scriptstyle(b)}}{{\leq}}\sum_{i=(t-1)(\tau-1)+1}^{t(\tau-1)}\sum_{j=t\cdot\tau+1}^{n}b^{j}d\left(\pi^{o}(i),\pi^{o}(i+1)\right)$
$\leq\sum_{i=(t-1)(\tau-1)+1}^{t(\tau-1)}\frac{b^{t\cdot\tau+1}}{1-b}d\left(\pi^{o}(i),\pi^{o}(i+1)\right)$
@@
where $d_{\max}\coloneqq\max_{u,v\in U}d\left(u,v\right)$. Inequality $(a)$ holds by the definition of $\ell_{\tau}$ and our algorithm, and inequality $(b)$ holds because $p_{O^{o}_{j}}\leq b^{j}$.
@@
(D.6) $\ell_{\tau}$ $\stackrel{{\scriptstyle(a)}}{{\geq}}\ell_{2}\stackrel{{\scriptstyle(b)}}{{\geq}}p_{\hat{O_{2}}}d\left(\hat{\pi}(1),\hat{\pi}(2)\right)\geq a^{2}d_{\max}.$
@@
Appendix E Ommited Experimental Settings and Results
E.1. Omitted Baseline
@@
$p_{i}^{-\alpha}+d(i,R)^{-\alpha}-1]^{-1/\alpha}.$
@@
E.2. More Experimental Results
We report the running time of our proposed algorithms and the baselines in Table 10.
Table 10. Run times of proposed algorithms and baselines (sec). The results marked with are obtained using the $\small\textsc{B}\tau\textsc{I-H}$ heuristic
Random EXPLORE DUM MSD MMR DPP B2I B3I B4I SB3I SB4I
Coat 0.83 2.81 0.97 1.08 1.24 0.96 1.41 4.70* 53.03* 6.13* 101.83*
KuaiRec 21.00 134.82 25.40 43.92 58.37 106.47 198.61 55.27* 283.62* 79.03* 294.47*
Netflix 10.76 180.60 14.51 14.71 14.51 50.67 181.57 72.79* 866.86* 72.61* 1322.30*
Movielens 40.12 787.52 68.91 135.89 87.50 561.19 929.23 154.21* 1121.35* 172.44* 1248.21*
Yahoo 53.78 1922.62 132.07 252.97 125.09 1316.29 2089.38 388.58* 3807.37* 399.16* 1448.03*
LETOR 0.17 9.90 2.71 1.94 1.74 2.12 0.93 10.05 873.70 10.07 772.58
LTRC 0.45 28.30 6.52 4.98 4.97 5.95 2.24 42.11 3584.67 36.15 3888.68